# Sample-efficient Unsupervised Policy Cloning from Ensemble Self-supervised Labeled Videos

### v1

@@ III-B1 Motivation
To predict expert actions accurately, the video labeling model $V(a^{v}_{i}|o^{v}_{i},o^{v}_{i+1})$ should be able to capture and understand which relative differences (between neighboring images) are similar and close, which requires it to understand the expert-related dynamics hidden in the videos. To this end, we introduce a latent future reconstruction task into model training, where another world model is employed. The world model takes the latent feature of the current observation and $V$-predicted latent action as inputs, trying to reconstruct the latent feature of the next observation. By supporting the world model to reconstruct different future observations, the video labeling model is forced to understand dynamics, refining more action-like information into a small-dimension output.
@@ III-B2 Implementation
The latent future reconstruction task updates the video labeling model and an extra world model jointly. First, a batch of observation pairs $\{(o^{v}_{i},o^{v}_{i+1})\}_{i=1}^{N}$ are sampled from expert video dataset $B^{v}$. Each current observation $o^{v}_{i}$ is encoded by the $f$ to obtain $f(o^{v}_{i})$ while each next observation $o^{v}_{i+1}$ is encoded by the $f$ to obtain $f(o^{v}_{i+1})$. Then, they are concatenated and further processed by latent predictor $g$ to produce the predicted latent action $z^{v}_{i}$. Then, the $f(o^{v}_{i})$ and $z^{v}_{i}$ are concatenated as the input to the world model $M_{e}(\cdot|z^{v}_{i},o^{v}_{i})$ and we can obtain a predicted observation $p^{v}_{i+1}$ in the latent space. Then we minimize the difference between the latent embeddings of predicted observation $p^{v}_{i+1}$and ground-truth next observation $f(o^{v}_{i+1})$, jointly updating the parameters of our video labeling model $V$ and world model $E_{w}$:
@@ III-C1 Motivation
If the video labeling model $V(a^{v}_{i}|o^{v}_{i},o^{v}_{i+1})$ can understand the expert video well, it should also generalize well to demonstrations sampled by other policies. To this end, we train our policy $\pi(a|o)$ to imitate the $V$-labeled videos through behavior cloning and utilize it to interact with environments for real transitions collection. Then $V$ is updated by achieving a real action prediction task on the collected real transitions. With the $V$ improved by all three tasks (visual shift contrast, latent future reconstruction, and ground-truth action prediction), the derived policy $\pi$ is also improved and in turn collects more action-labeled data of higher quality for the ground-truth action prediction. This task is not only set to provide the $V$ with the information of real action spaces, but also serves as a generalization task to improve both the robustness and performance of $V$ by different data distribution and optimization targets.
@@ III-C2 Implementation
First, a batch of transitions $\{(o^{e}_{i},a^{e}_{i},o^{e}_{i+1})\}^{N}_{i=1}$ are sampled from environmental transitions $B^{e}$. We feed each observation pair into our video labeling model $V$ and obtain the $V$-predicted action $a^{e,v}_{i}$. The difference between predicted action vector $a^{e,v}_{i}$ and ground-truth action $a^{e}_{i}$ is minimized. The objective is formulated as the following:
$\mathcal{L}_{GAP}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{C}(a^{e,v}_{i,(c)}loga^{e}_{i,(c)}).$ (5)
where the subscript $(c)$ denotes the $c$-th dimension of a vector.
TABLE I: The few-shot policy learning performance comparison between UPESV and all baseline methods. Each method are only allowed 100K steps on each task, which require methods to have high sample efficiency to obtain an effective policy. The proposed UPESV exhibits state-of-the-art performance when compared with both advanced LFV methods and RL methods.
@@
Bigfish 30.5 ± 1.6 3.6 ± 3.7 0.9 ± 0.0 20.6 ± 0.7 0.9 ± 0.1 1.0 ± 0.1
Maze 9.7 ± 0.2 7.4 ± 2.4 0.8 ± 1.5 9.6 ± 0.1 5.0 ± 0.7 4.6 ± 0.7
Heist 9.4 ± 0.3 7.6 ± 1.9 1.2 ± 1.2 9.4 ± 0.3 3.7 ± 0.2 3.4 ± 0.9
Coinrun 7.4 ± 0.2 6.7 ± 0.9 2.4 ± 0.2 6.2 ± 0.4 4.1 ± 0.5 3.6 ± 1.0
Plunder 3.5 ± 0.7 4.2 ± 0.3 4.0 ± 0.7 4.8 ± 0.1 4.4 ± 0.4 4.5 ± 0.4
Dodgeball 9.1 ± 0.8 5.4 ± 1.1 0.0 ± 0.1 5.9 ± 1.1 1.1 ± 0.2 0.6 ± 0.2
Jumper 6.6 ± 0.2 6.4 ± 0.3 0.6 ± 0.4 7.3 ± 0.2 3.5 ± 0.7 3.2 ± 0.5
Climber 6.8 ± 0.6 3.3 ± 0.2 0.1 ± 0.2 4.7 ± 0.3 2.2 ± 0.2 2.4 ± 0.5
Fruitbot 20.6 ± 1.6 3.6 ± 0.1 -1.4 ± 0.2 0.5 ± 0.3 -1.9 ± 1.0 -2.4 ± 0.9
Starpilot 15.0 ± 0.8 12.8 ± 13.9 2.1 ± 0.1 20.3 ± 1.6 2.6 ± 0.9 1.8 ± 0.6
Ninja 6.3 ± 0.3 4.2 ± 1.1 0.3 ± 0.5 5.2 ± 0.1 3.4 ± 0.3 3.0 ± 0.3
Miner 9.3 ± 1.2 5.8 ± 1.3 0.7 ± 0.3 6.7 ± 0.6 1.2 ± 0.2 1.2 ± 0.2
Caveflyer 3.5 ± 0.6 2.8 ± 1.1 2.7 ± 1.1 3.9 ± 0.1 3.0 ± 0.4 3.2 ± 0.6
Leaper 2.9 ± 0.3 2.5 ± 0.5 0.2 ± 0.8 2.7 ± 0.2 2.6 ± 0.3 2.5 ± 0.4
Chaser 0.8 ± 0.1 0.8 ± 0.0 0.1 ± 0.1 0.8 ± 0.0 0.4 ± 0.2 0.6 ± 0.2
Bossfight 2.0 ± 0.4 0.4 ± 0.3 0.0 ± 0.0 0.3 ± 0.3 0.1 ± 0.1 0.2 ± 0.2
Mean score 9.0 4.8 0.9 6.8 2.3 2.1
@@ III-D Unsupervised Policy Cloning
As mentioned in Section III.C, our policy $\pi(a|o)$ and video labeling model $V(a^{v}_{i}|o^{v}_{i},o^{v}_{i+1})$ are trained and improved jointly. $\pi(a|o)$ imitates the $V$-labeled expert videos through behavior cloning and interacts with environments for self-supervised data collection, while $V$ is improved for a better $\pi(a|o)$ through three self-supervised tasks ($\mathcal{L}_{sm}$, $\mathcal{L}_{lfr}$, and $\mathcal{L}_{sm}$) conducted over the expert videos and collected transitions. After a sample-efficient iterative training process, an effective $\pi(a|o)$ and an advanced labeling model $V$ are both obtained.
The behavior cloning process of $\pi(a|o)$ is described here. For every observation pairs $(o^{v}_{i},o^{v}_{i+1})$ from $B^{v}$, we use the video labeling model $V$ to obtain a action vector $a^{v}_{i}$. Then the action-free observation pairs are labeled with $a^{v}_{i}$, denoted as $(o^{v}_{i},a^{v}_{i},o^{v}_{i+1})$. Based on the current observation $o^{v}_{i}$ and the policy $\pi(a|o)$, we obtain the policy-based action vector $a^{v,\pi}_{i}$ and then minimize its difference from label $a^{v}_{i}$ for policy learning:
$\mathcal{L}_{UPC}=\frac{1}{N}\sum_{i=1}^{N}||a^{v,\pi}_{i}-a^{v}_{i}||^{2}.$ (6)
@@ IV-A Experimental Settings
We evaluated our method on all sixteen environments introduced by the Procgen benchmark [46]. Procgen provides a diverse set of procedurally-generated video game environments, each with unique challenges, multiple dynamics, and changing visual style. For example, in Maze task, the screen color, map size, and maze content are all changing drastically, which places high demands on both representation learning and dynamics understanding of methods.
@@
We compare our UPESV with five current advanced online policy learning baselines, including both Learning From Video (LFV) methods and pure RL methods: BCO [21], ILPO [22], LAPO [43], PPO [44], and TVL [45]. Among them, the first two methods are current state-of-the-art Imitation Learning from only Observation (ILO) methods, requiring no other supervisions like UPESV. Our UPESV differs from these ILO baselines apparently. They either train a video labeling model for behavior cloning only on environmental transitions (BCO) or extract environmental dynamics from only expert videos (ILPO). By contrast, UPESV takes advantages from both kinds of methods, employing two different self-supervised tasks to utilize both datasets simultaneously for an advanced video labeling model. In addition, the additional visual self-supervised task further improves the representation ability and performance of UPESV. LAPO is a state-of-the-art video pre-training method. The pre-trained latent policy is finetuned by rewards and actions in environmental interactions. Inverse RL methods also provide solutions to learn video-based policy online without other supervisions, but their goal is to restore the expert reward as much as possible, which makes them difficult to surpass RL methods exposed to expert rewards [38, 39], in terms of sample efficiency [20]. Considering that they are generally not evaluated on Procgen, we directly employ TVL and PPO, two state-of-the-art RL methods on Procgen provided with expert rewards in comparison. Note that our UPESV doesn’t access these expert rewards and any extra supervisions.
@@
We choose full distributions of levels and easy mode in Procgen throughout this paper. For each method, 100K steps of environmental interactions are allowed. For methods requiring expert videos, 60K update times (during 100K environmental steps) are allowed for the training of each model (including policy) in their frameworks. The expert videos containing 8M steps are generated by a well-trained RL-based policy, which is provided by previous work [43]. For methods requiring rewards, we allow them to access the expert reward provided by Procgen environments. The architectures of UPESV follow previous works [44, 43, 45] that are currently advanced on Procgen. Adam [47] is chosen as the optimizer throughout the paper, with the batch size set to 128. The learning rate is set to 3e-5 for the visual shift contrast task, 3e-4 for the latent future reconstruction task, 1e-3 for the ground-truth action prediction task, and 2e-4 for unsupervised policy cloning. The maximum shift distance $s$ in visual shift contrast task is set to 1, which performs better than others in experiments. The results of UPESV are over average 5 random seeds.
@@
Figure 3: Ablation study on all three self-supervised tasks. Each task is employed with reasonable motivation necessary in UPESV.
@@ IV-B Sample Efficiency Comparison
We test the policy learning abilities of all the methods when interaction is limited, which directly reflects their sample efficiency. Each method is allowed to interact with environments for only 100K steps. 100K steps are much less than the 25M employed in previous works [46, 45], requiring extremely high sample efficiency. The results across all 16 tasks are shown in Table 1. Compared with three video-based baselines, UPESV performs better on 12/16 tasks and not worse on the others except Starpilot. These advantages demonstrate the UPESV video labeling model has a better understanding of both visual inputs and environmental dynamics, which can be attributed to the employed three self-supervised tasks. We observe trivial solutions that lead to huge performance drops in ILPO. This is due to its mode collapse in complex environments, which is also observed by [20, 43]. Compared with two advanced RL methods, UPESV’s performance advantages are very obvious. In addition, video-based methods BCO and LAPO also overall perform better than pure RL methods. Note that both UPESV and BCO are reward-free. These phenomena demonstrate that the videos may serve as a more efficient teacher than expert rewards, especially when interactions are limited. In summary, the huge efficiency advantages enable UPESV to have state-of-the-art few-shot policy learning ability.
@@ IV-C Ablations: Necessity of Each Self-supervised Task
In UPESV, we employ three different self-supervised tasks to jointly learn the video labeling model and policy. These three tasks are organically combined, where each one performs its own duties. To demonstrate that they are all indispensable in our framework, we sequentially ablate the latent future reconstruction task (UPESV w/o LFR), ground-truth action prediction task (UPESV w/o GAP), and visual shift contrast task (UPESV w/o VSC), observing the performance changes. In UPESV w/o GAP, we have to train the video labeling model without accessing true action space, so we train a policy decoder to remap the latent actions to real actions based on interactions after the policy learning process, which follows previous works [22]. The results are shown in Fig. 3, demonstrating that (i) every employed self-supervised task is necessary in UPESV, and (ii) the visual shift contrast task is crucial for the model to benefit from the other two tasks for advanced environment dynamics understanding.

### v2

@@ III-B1 Motivation
To predict expert actions accurately, the video labeling model $V(a^{v}_{i}|o^{v}_{i},o^{v}_{i+1})$ should be able to capture and understand which relative differences (between neighboring images) are similar and close, which requires it to understand the expert-level environmental dynamics hidden in the videos. To this end, we introduce a latent future reconstruction task into model training, where another world model is employed. The world model takes the latent feature of the current observation and $V$-predicted latent action as inputs, trying to reconstruct the latent feature of the next observation. By supporting the world model to reconstruct different future observations in the videos, the labeling model is forced to understand the dynamics and refine more useful information into its small-dimension outputs.
@@ III-B2 Implementation
The latent future reconstruction task updates the video labeling model $V$ and an extra world model $G_{w}$ jointly. First, a batch of observation pairs $\{(o^{v}_{i},o^{v}_{i+1})\}_{i=1}^{N}$ is sampled from the expert video dataset $B^{v}$. Each current observation $o^{v}_{i}$ is encoded by the $f$ to obtain $f(o^{v}_{i})$, while each next observation $o^{v}_{i+1}$ is encoded by the $f$ to obtain $f(o^{v}_{i+1})$. Then, they are concatenated and further processed by latent predictor $g$ to produce the predicted latent action $z^{v}_{i}$. $z^{v}_{i}$ is discretized by vector quantization [54] (the variable name is not changed for simplicity). Then, the $f(o^{v}_{i})$ and $z^{v}_{i}$ are concatenated as the input to the world model $G_{w}(\cdot|z^{v}_{i},f(o^{v}_{i}))$, and we can obtain a predicted observation $p^{v}_{i+1}$ in the latent space. Then we minimize the difference between $p^{v}_{i+1}$ and the latent embeddings of the ground-truth next observation $f(o^{v}_{i+1})$, jointly updating the parameters of our video labeling model $V$ and the world model $G_{w}$. The gradient is stopped before $f$ in backpropagation. The learning objective of this task is defined as the following:
@@ III-C1 Motivation
If the video labeling model $V(a^{v}_{i}|o^{v}_{i},o^{v}_{i+1})$ can understand the expert video well, it should also generalize well to demonstrations sampled by other policies. To this end, we train our policy $\pi(a|o)$ to imitate the $V$-labeled videos through behavior cloning and utilize it to interact with environments for real transitions collection. Then $V$ is updated by achieving a real action prediction task on the collected real transitions (random collection in the first round). With the $V$ improved by learning the real-action-labeled transitions, the derived policy $\pi$ is also improved and in turn collects more action-labeled data of higher quality for the ground-truth action prediction. This task not only provides the $V$ with the information of real action spaces, but also serves as a generalization task to improve both the robustness and performance of $V$ by different data distribution and optimization targets. Since the other two self-supervised tasks don’t require real transitions, they can be carried out before this task and environmental interactions to further improve the sample efficiency.
@@ III-C2 Implementation
First, a batch of transitions $\{(o^{e}_{i},a^{e}_{i},o^{e}_{i+1})\}^{M}_{i=1}$ is sampled from the environmental transition buffer $B^{e}$. We feed each observation pair into our video labeling model $V$ and obtain the $V$-predicted action $a^{e,V}_{i}$. The difference between the predicted action vector $a^{e,V}_{i}$ and the ground-truth action $a^{e}_{i}$ (one-hot) is minimized. The gradient is stopped before $g$ to reduce burden. The learning objective of this task is formulated as the following:
$\mathcal{L}_{GAP}=-\frac{1}{M}\sum_{i=1}^{M}\sum_{c=1}^{C}(a^{e,V}_{i,(c)}loga^{e}_{i,(c)}),$ (5)
where the subscript $(c)$ denotes the $c$-th vector dimension.
TABLE I: The interaction-limited policy learning performance comparison between UPESV and all baseline methods. Each method is only allowed 100k steps on each task, which requires methods to have high sample efficiency to obtain an effective policy. The proposed UPESV exhibits state-of-the-art performance when compared with both advanced LFV methods and RL methods.
@@
Bigfish 30.5 ± 1.6 3.6 ± 3.7 0.8 ± 0.1 20.6 ± 0.7 0.9 ± 0.1 1.0 ± 0.1
Maze 9.7 ± 0.2 7.4 ± 2.4 4.2 ± 0.3 9.6 ± 0.1 5.0 ± 0.7 4.6 ± 0.7
Heist 9.4 ± 0.3 7.6 ± 1.9 6.7 ± 0.5 9.4 ± 0.3 3.7 ± 0.2 3.4 ± 0.9
Coinrun 7.4 ± 0.2 6.7 ± 0.9 3.7 ± 1.3 6.2 ± 0.4 4.1 ± 0.5 3.6 ± 1.0
Plunder 3.5 ± 0.7 4.2 ± 0.3 2.2 ± 1.3 4.8 ± 0.1 4.4 ± 0.4 4.5 ± 0.4
Dodgeball 9.1 ± 0.8 5.4 ± 1.1 0.6 ± 0.1 5.9 ± 1.1 1.1 ± 0.2 0.6 ± 0.2
Jumper 6.6 ± 0.2 6.4 ± 0.3 3.1 ± 0.6 7.3 ± 0.2 3.5 ± 0.7 3.2 ± 0.5
Climber 6.8 ± 0.6 3.3 ± 0.2 3.4 ± 0.5 4.7 ± 0.3 2.2 ± 0.2 2.4 ± 0.5
Fruitbot 20.6 ± 1.6 3.5 ± 0.5 -2.0 ± 0.7 0.5 ± 0.3 -1.9 ± 1.0 -2.4 ± 0.9
Starpilot 15.0 ± 0.8 12.8 ± 13.9 0.5 ± 0.7 20.3 ± 1.6 2.6 ± 0.9 1.8 ± 0.6
Ninja 6.3 ± 0.3 4.2 ± 1.1 2.2 ± 1.1 5.2 ± 0.1 3.4 ± 0.3 3.0 ± 0.3
Miner 9.3 ± 1.2 5.8 ± 1.3 1.2 ± 0.4 6.7 ± 0.6 1.2 ± 0.2 1.2 ± 0.2
Caveflyer 3.5 ± 0.6 2.8 ± 1.1 3.2 ± 0.3 3.9 ± 0.1 3.0 ± 0.4 3.2 ± 0.6
Leaper 2.9 ± 0.3 2.5 ± 0.5 2.6 ± 0.2 2.7 ± 0.2 2.6 ± 0.3 2.5 ± 0.4
Chaser 0.8 ± 0.1 0.8 ± 0.0 0.7 ± 0.0 0.8 ± 0.0 0.4 ± 0.2 0.6 ± 0.2
Bossfight 2.0 ± 0.4 0.4 ± 0.3 0.1 ± 0.0 0.3 ± 0.3 0.1 ± 0.1 0.2 ± 0.2
Mean score 9.0 4.8 2.1 6.8 2.3 2.1
@@ III-D Unsupervised Policy Cloning
As mentioned in Section III.C, our policy $\pi(a|o)$ and video labeling model $V(a^{v}_{i}|o^{v}_{i},o^{v}_{i+1})$ are improved jointly. $\pi(a|o)$ imitates the $V$-labeled expert videos through behavior cloning and interacts with environments for self-supervised data collection. With $V$ improved by the $\pi$-collected training data, $\pi(a|o)$ is also further enhanced and in turn provides more and better self-supervised training data. After a sample-efficient and iterative training process, an effective $\pi(a|o)$ and an advanced video labeling model $V$ are both obtained.
$\pi$ consists of three networks: a feature encoder $f^{\pi}$, a policy predictor $g^{\pi}$, and an action projector $h^{\pi}$. To reduce the training burden, we share the visual encoder and action projector of the labeling model $V$ with those of $\pi$, i.e., $f^{\pi}=f$ and $h^{\pi}=h$, only training $g^{\pi}$ here. For higher sampling and training efficiency, $g^{\pi}$ is trained after learning $f^{\pi}$ (i.e., $f$) on videos and before learning $h^{\pi}$ (i.e., $h$) with interactions. For every observation pair $(o^{v}_{i},o^{v}_{i+1})$ from the video dataset $B^{v}$, we use $V$ to obtain the predicted latent action $z^{v}_{i}$. Based on the current observation $o^{v}_{i}$ and the policy $\pi$, we obtain the policy-derived latent action $z^{v,\pi}_{i}=g^{\pi}(f^{\pi}(o^{v}_{i}))$. Then we minimize the difference between $z^{v}_{i}$ and $z^{v,\pi}_{i}$:
$\mathcal{L}_{UPC}=\frac{1}{N}\sum_{i=1}^{N}||z^{v,\pi}_{i}-z^{v}_{i}||^{2}.$ (6)
The gradient is stopped before $f^{\pi}$. The policy-derived real action $a^{v,\pi}_{i}$ can be obtained by $a^{v,\pi}_{i}=h^{\pi}(z^{v,\pi}_{i})$.
@@ IV-A Experimental Settings
We evaluated our method on all sixteen environments introduced by the Procgen benchmark [58]. Procgen provides a diverse set of procedurally generated video game environments, each with unique challenges, multiple dynamics, and changing visual styles. For example, in Maze task, the screen color, map size, and maze content are all changing drastically, which places high demands on both representation learning and dynamics understanding of learning methods.
@@
We compare our UPESV with five current advanced online policy learning baselines, including both Learning From Video (LFV) methods and pure RL methods: BCO [28], ILPO [29], LAPO [55], PPO [56], and TVL [57]. Among them, the first two methods are current state-of-the-art Imitation Learning from only Observation (ILO) methods, requiring no other supervision like UPESV. They either train a video labeling model for behavior cloning only on environmental transitions (BCO) or extract environmental dynamics from only expert videos (ILPO), while UPESV differs from these ILO baselines, employing several self-supervised tasks to utilize both datasets simultaneously for an advanced video labeling model. LAPO is a state-of-the-art video pre-training method. The pre-trained latent policy is finetuned by rewards and actions in environmental interactions. Inverse RL methods also provide solutions to learn video-based policy online without other supervision, but their goal is to restore the expert reward as much as possible, which makes them difficult to surpass RL methods exposed to expert rewards [47, 48], in terms of sample efficiency [27]. Considering that they are generally not evaluated on Procgen, we directly employ TVL and PPO, two state-of-the-art RL methods on Procgen provided with expert rewards in comparison. Note that our UPESV doesn’t access these expert rewards and any extra supervision.
@@
We choose full distributions of levels and easy mode in Procgen throughout this paper. For each method, 100k steps of environmental interactions are allowed. The expert videos required by the video-based methods contain 8M steps. They are generated by well-trained RL-based policies, provided by [55]. For methods requiring rewards, we allow them to access the expert reward provided by Procgen environments. UPESV employs IMPALA-CNN [59] as encoders and predictors, MLP as projectors, and U-NET [60] as the world model. These follow previous works [56, 55, 57] that are currently advanced on Procgen. The number of parallel environments is $64$, and the update frequency is $64$. The update times are 3k for ground-truth action prediction, 50k for unsupervised policy cloning, and 60k for the other objectives. Adam [61] is chosen as the optimizer, with the batch size set to $128$ for expert videos and $512$ for environmental transitions. The learning rate is set to $53e-$ 5 for the visual shift contrast, $43e-$ 4 for the latent future reconstruction, $31e-$ 3 for the ground-truth action prediction, and $42e-$ 4 for unsupervised policy cloning. The maximum shift distance $s$ in visual shift contrast is $1$, and the EMA momentum is set to $0.05$. The results of UPESV are obtained through $5$ different runs.
@@
Figure 3: Ablation study of three self-supervised tasks on eight procgen environments. Each task is necessary in the proposed UPESV.
@@ IV-B Sample Efficiency Comparison
We test the policy learning abilities of all the methods when interaction is limited, which directly reflects their sample efficiency. Each method is allowed to interact with environments for only 100k steps. 100k steps are much less than the 25M employed in previous works [58, 57], requiring extremely high sample efficiency. The results across all 16 tasks are shown in Table 1. Compared with three video-based baselines, UPESV performs better on 12/16 tasks and is also competitive on the others. These advantages demonstrate the UPESV video labeling model has a better understanding of both visual inputs and environmental dynamics, which can be attributed to the employed three self-supervised tasks. We observe trivial solutions that lead to performance drops in ILPO. This is due to its mode collapse in complex environments, which is also observed by [27, 55]. Compared with two advanced RL methods, UPESV’s performance advantages are very obvious. In addition, video-based methods BCO and LAPO also overall perform better than pure RL methods. Note that both UPESV and BCO are reward-free. These phenomena demonstrate that the videos may serve as a more efficient teacher than expert rewards, especially when interactions are extremely limited. In summary, UPESV exhibits advanced sample efficiency without exposure to expert rewards or actions.
@@ IV-C Ablations: Necessity of Each Self-supervised Task
In UPESV, we employ three different self-supervised tasks to jointly learn the video labeling model and policy. These three tasks are organically combined, where each one performs its own duties. To demonstrate that they are all indispensable in our framework, we sequentially ablate the latent future reconstruction task (UPESV w/o LFR), the ground-truth action prediction task (UPESV w/o GAP), and the visual shift contrast task (UPESV w/o VSC), observing the performance changes. In UPESV w/o GAP, we have to train the video labeling model and clone policies without accessing the true action space. To this end, we additionally train an action decoder for the cloned policy to remap the predicted latent actions to real actions based on environmental interactions and the world model, similar to [29]. The results are shown in Fig. 3, demonstrating that each employed self-supervised task is necessary for the proposed UPESV to achieve ideal policy cloning performance.